a. According to the Rule of 70, how many years will it take a country to double its output at each of the following annual growth rates?

0.5 percent: ___ years

1 percent: ____ years

1.4 percent: ____ years

2 percent: ____ years

2.8 percent: ___ years

3.5 percent: ____ years

7 percent: ____ years
1
Expert's answer
2014-11-19T13:13:22-0500
According to the Rule of 70, to find how many years will it take a country to double its output at each of the following annual growth rates 70 is divided by the interest percentage per period. So:
0.5 percent: 70/0.5 = 140 years
1 percent: 70/1 = 70 years
1.4 percent: 70/1.4 = 50 years
2 percent: 35 years
2.8 percent: 25 years
3.5 percent: 20 years
7 percent: 10 years.